# Understanding the Skill Gap in Recurrent Language Models: The Role of the Gather-and-Aggregate Mechanism

### v1

@@ 3 A Case Study: Retrieval in MMLU
MMLU covers 57 tasks that include elementary mathematics, US history, computer science, law, and more. To achieve high accuracy in this test, models must have extensive world knowledge and problem solving ability (Hendrycks et al., 2021). MMLU covers 57 diverse tasks, including elementary mathematics, U.S. history, computer science, and law. Thus, models must possess extensive world knowledge to perform well (Hendrycks et al., 2021). In addition to the knowledge requirement, the MMLU evaluation format requires selecting a letter label (A-D) and retrieving it rather than scoring complete answers, demanding a basic form of retrieval ability.
@@ D. Access Port
SSMs tend to perform worse on MMLU unless trained significantly longer (Waleffe et al., 2024), despite matching Transformer performance on many knowledge benchmarks. Given known retrieval limitations of SSM architectures (Jelassi et al., 2024), we ask whether their performance gap with Transformers on MMLU stems from difficulties in retrieving rather than acquisition of knowledge?
This section presents two key findings:
@@
    Models can perform well on MMLU with limited knowledge, suggesting that, for the evaluated models, the task primarily reflects in-context retrieval. This adds to concerns raised by prior work, including Wang et al. (2024b); Zhao et al. (2024).
@@
These insights, combined with earlier findings about perplexity gaps between attention-based models and SSM variants on the Pile dataset (Arora et al., 2023), support our main thesis: the general performance gap between Transformer- and SSM-based language models can be traced to specific algorithmic capabilities within a small subset of heads.
Our investigation compares the attention-based Transformer, Llama-3.1-8B-Instruct, with two SSM-based architectures: Falcon-Mamba-7B-Instruct (based on Mamba-1) and Llamba-8B-untied (based on Mamba-2). For brevity, we omit instructional and embedding-tying descriptors throughout our analysis. We evaluate performance on MMLU alongside a suite we term Knowledge-Focused Tasks. These include ARC-Challenge, ARC-Easy, PIQA, Winogrande, OpenBookQA, and HellaSwag—tasks selected because they primarily assess factual knowledge and commonsense reasoning with minimal dependence on retrieval from recent context. Our analysis follows a bottom-up approach, identifying discrepancies between MMLU and other knowledge-focused tasks, then tracing them back to retrieval. Specifically, we demonstrate:
  • Section 3.1 outlines the benchmarks used in our analysis, distinguishing MMLU from knowledge-focused tasks based on their reliance on retrieval.
  • Section 3.2 shows that MMLU performance initially appears to depend on a specific critical layer rather than the network’s overall stored knowledge, based on cumulative layer pruning.
  • Section 3.3 builds on this finding through more targeted experiments, revealing that MMLU actually relies on the coordination between two essential layers, both of which must be present for successful performance.
  • Section 3.4 refines this perspective by shifting from layers to individual heads, revealing that MMLU disproportionately relies on a few specific heads rather than distributed knowledge.
  • Section 3.5 demonstrates these heads are part of a retrieval mechanism consistently present across all architectures.
Figure 3: Performance of Llama-3.1-8B-Minimal with layers removed and restored. We remove one layer at a time, evaluate performance on MMLU and knowledge tasks’ score, and then restore the layer before proceeding to the next. The results reveal that both layers 16 and 17 are critical for MMLU performance, as pruning either of the last two layers of the minimal model causes a significant drop on MMLU scores while causing only a minor drop in scores on knowledge-focused tasks.
@@
Throughout the paper, we evaluate performance on two categories of benchmarks: MMLU and what we term Knowledge-Focused Tasks. MMLU’s format distinctively presents multiple-choice questions across 57 diverse subjects, requiring models to process lengthy contexts containing both questions and answer choices, then identify the correct option by outputting the corresponding letter label (typically A, B, C, D). While MMLU is widely used as a benchmark, some researchers have raised questions about its evaluation approach (Alzahrani et al., 2024; Wang et al., 2024b; Zhao et al., 2024) For instance, Wang et al. (2024b) argue that MMLU performance may not necessarily indicate true comprehension by large language models, and studies by Zhao et al. (2024) have highlighted issues of benchmark contamination that inevitably affect evaluation integrity.
In contrast, our Knowledge-Focused Tasks—including ARC-Challenge and ARC-Easy (Clark et al., 2018), PIQA (Bisk et al., 2019), Winogrande (Sakaguchi et al., 2019), OpenBookQA (Mihaylov et al., 2018), and HellaSwag (Zellers et al., 2019)—were selected because they primarily assess factual knowledge and commonsense reasoning with minimal dependence on retrieval from recent context. The fundamental difference lies in what the models are asked to predict: while MMLU requires models to retrieve a letter label (A, B, C, or D) after processing question-answer pairs, these other benchmarks are evaluated by scoring the likelihood each model assigns to full candidate answers. In this setup, we retrieve the complete text of each answer option in advance and present it directly to the model. The model’s task is then to assign a likelihood score to each one, rather than selecting a label. This evaluation approach removes the burden of interpreting or extracting answer choices from the context and instead emphasizes how well the model’s internal knowledge supports each complete answer.
@@ 3.2 A Crucial Layer for MMLU
Motivated by the struggle of SSMs with MMLU (Waleffe et al., 2024), and by the fact that Transformers use specialized components to solve multiple-choice questions (Lieberum et al., 2023), we examine whether solving this task utilizes different components in SSM-based language models as well. We start our investigation by isolating specific layers’ contributions. Since all language models utilize residual skip connections, eliminating a layer could reveal its specific contribution, while keeping the rest of the information flowing.
We iteratively remove one layer at a time from the model’s end and measure performance on MMLU against knowledge-focused tasks. For Llama-3.1-8B and Llamba-8B, we remove entire layers, including both mixer and MLP components, since MLPs likely depend on mixer outputs. The experiments provided in Figure 9 show similar results when the MLPs are kept intact.
Figure 2 presents the performance trends for Llama-3.1-8B on both knowledge tasks and MMLU as layers are progressively removed, and Figure 8 extends this analysis to Falcon-Mamba-7B and Llamba-8B models. All architectures revealed a striking pattern: while performance on knowledge-focused benchmarks decreased uniformly as layers were pruned, MMLU performance remained remarkably stable until a specific layer was removed. At this point, MMLU performance dropped dramatically, suggesting this particular layer encodes crucial skills required for MMLU. This observation also raises an intriguing question about how these models can maintain high MMLU scores while apparently retaining only a fraction of their knowledge capacity.
@@ 3.3 MMLU Relies on Two Crucial Layers
Based on insights from Figure 2, we refer to the minimal model as the smallest subnetwork that retains all layers up to (and including) the layer before the sharp drop in MMLU performance. We now repeat the previous experiment, this time on the minimal model, with a slight modification. Instead of cumulatively removing one layer at a time from the end of the network, we remove a single layer, evaluate the model performance without it, and then restore the layer before proceeding to the next.
Figure 3 illustrates the results of this experiment in Llama-3.1-8B, and Figure 10 shows the results for Falcon-Mamba-7B and Llamba-8B as well. The findings reveal a surprising insight on the minimal model: not one, but two layers are critical for the performance of MMLU. Unlike earlier layers, the last two layers of the minimal model show no correlation between MMLU score and knowledge task performance. Pruning either layer causes a significant drop in MMLU accuracy while causing only a minor drop in scores on knowledge-focused tasks, underscoring their essential role in the task.
@@
Table 1: Evaluating the Role of Critical Heads in Llama-3.1-8B-Minimal’s Performance. This table presents the impact of two critical attention heads (L16H22 and L17H25) on MMLU performance while keeping all preceding layers intact. Four configurations were tested: (1) both heads retained, (2) only the last-layer head retained, (3) only the second-to-last-layer head retained, and (4) neither head retained. Results show that these heads must work together to sustain high MMLU performance ($>$66%), while removing either or both drops accuracy to random guessing (25%).
@@ 3.4 Zooming In: MMLU Relies on Two Crucial Heads
Having established the importance of two specific layers in Section 3.3, we now refine our analysis from the layer level to the head level, investigating how individual attention heads within these critical layers contribute to each language model’s performance.
To precisely identify the crucial heads, we conducted a systematic ablation study on the minimal model defined in Section 3.3 (where only the last two layers are critical). We systematically zeroed out the outputs of individual heads in these two critical layers and measured the impact on both MMLU and knowledge-focused tasks. This analysis revealed a remarkably small subset of heads whose removal drastically reduced MMLU accuracy while leaving knowledge benchmark performance intact. We designated these as ”critical heads” - specifically, the heads whose removal significantly impacted MMLU scores.
To further validate the importance of these identified critical heads, we removed all noncritical heads from the last two layers and evaluated performance across four configurations: (1) no critical heads retained, (2) only critical heads from the last layer retained, (3) only critical heads from the second last layer retained, and (4) critical heads in both layers retained. The results, presented in Table 1 for Llama-3.1-8B and in Tables 6 and 5 for the other models, demonstrate that MMLU performance depends on just a few mixing heads rather than the model’s broader knowledge representation. Disabling these specific heads causes accuracy to plummet to random guessing levels. Remarkably, in Llama-3.1-8B, a single head is sufficient to achieve over 66% MMLU accuracy, while in Falcon-Mamba-8B, just four Mamba-1 channels reach 52.26% accuracy.
@@ 3.5 Retrieval Emerges as a Key Factor in MMLU
Our earlier findings demonstrated that two attention heads from distinct layers collaborate to effectively solve MMLU across all tested architectures. We now present evidence that these heads constitute part of a broader in-context retrieval mechanism. To establish this retrieval function, we analyze how removing one critical head impacts performance on a synthetic key-value (KV) retrieval task.
In the KV-Retrieval task, the model must recall a value associated with a given key from a predefined dictionary, testing its ability to store and access structured key-value pairs. For instance, given the prompt:
@@
We conducted experiments across 55 KV-Retrieval configurations with progressively increasing numbers of key-value pairs, each comprising 1,000 synthetic samples. For each configuration, we evaluated Llama-3.1-8B twice: once as-is and once with L17H24 (the head previously identified as critical in Section 3.4) removed. The results illustrated in Figure 4 reveal that removing L16H22 or L17H24 causes a decline in retrieval accuracy beyond just MMLU performance, confirming these heads’ essential role in the model’s retrieval mechanism. Notably, the impact of removing L17H24 grows with the number of key-value pairs, suggesting that its contribution to retrieval becomes more important as structural complexity increases. These findings extend previous research: the retrieval heads we identified align with and generalize the Content Gatherer head described by Lieberum et al. (2023) and the retrieval mechanisms documented by Wu et al. (2024). Our work demonstrates that similar retrieval mechanisms also emerge within SSM-based language models.
The pronounced role of retrieval in MMLU performance suggests that a broader Transformer-SSM performance gap may primarily stem from retrieval capabilities concentrated in just a few specific heads, rather than fundamental differences in language modeling capabilities. This insight fundamentally reframes our understanding of architectural tradeoffs, as the performance disparity appears to stem from a highly localized functional difference rather than a distributed architectural advantage.
In the next section, we will explore these two types of head and their distinct properties, revealing how they serve different roles.
@@ 4 Gather-and-Aggregate Heads
In Section 3, we identified critical heads responsible for in-context retrieval across both Transformer- and SSM-based language models. We now examine a shared computational pattern underlying these heads: the Gather-and-Aggregate (G&A) mechanism.
The mechanism unfolds in two stages. First, a Gather Head isolates meaningful segments within the input—such as multiple-choice answers—and summarizes each into a representative token. Then, a later Aggregate Head uses these summaries to retrieve the information most relevant to the prompt.
While previous work has described retrieval-oriented heads in Transformers—such as the “Correct Letter” heads identified by Lieberum et al. (2023)—these analyses focused narrowly on multiple-choice tasks. Our findings show that such heads are part of a broader retrieval strategy that generalizes beyond this format. This perspective also helps refine the interpretation of results from Wu et al. (2024), who treated a group of retrieval heads as playing the same role. We find that these heads instead specialize in complementary functions within a coordinated process.
Finally, we show that this same mechanism emerges in SSM-based models, despite their lack of attention. These models exhibit smoother and less discrete patterns, but still learn to implement a G&A strategy using different architectural tools.
The remainder of this section unpacks the mechanism in detail. We begin with how it operates in Transformers (Sections 4.1 and 4.2), then present experimental evidence that it also underlies retrieval in SSMs (Section 4.3).
@@ 4.1 Gather Heads in Transformers
A Gather Head compresses semantically related tokens within a segment into a single representative token. In Transformer models, this typically occurs through attention patterns that cause the final token of each segment to “attend to” all prior tokens in the same segment, effectively summarizing them into a compact representation (see Figure 1(a)).
For example, consider the multiple-choice format introduced in Section 3:
@@
While effective, this process is not always clean: attention sometimes leaks to unrelated tokens—particularly “attention sinks” Xiao et al. (2024) used to stabilize processing over long contexts. We return to these imperfections in Section 5.
@@ 4.2 Aggregate Heads in Transformers
Once the Gather Head condenses segments into representative tokens, an Aggregate Head—in a subsequent layer—retrieves and combines these summaries to inform the model’s prediction.
This head typically attends from a target token (e.g., the final “Answer:” position) over all gathered summaries, assigning greater weights to the most relevant one (see Figure 1(b)). This yields a sparse linear combination over segment summaries, often resembling an argmax-like operation (Tulchinskii et al., 2024):
@@ 4.3 G&A in Transformers and SSMs: Experimental Evidence
The G&A pattern is well-documented in Transformers. But do SSM-based models—like Mamba-1 and Mamba-2—implement the same structure? We answer this question through a two-part experimental procedure designed to isolate and test the role of specific heads:
  1. 1.
    Pruning: We prune each model to a minimal configuration (see Section 3.3), zeroing out all heads in the final two layers except those suspected to implement G&A (see Tables 6 and 1).
  2. 2.
    Masking: At runtime, we constrain the preserved heads’ attention patterns:
    • Gather Heads may only attend within each segment and to its final token.
    • Aggregate Heads may only attend to those final summary tokens.
    (See Figure 6 for an example.)
Across all models—Llama-3.1-8B (Transformer), Llamba-8B (Mamba-2), and Falcon-Mamba-7B (Mamba-1)—the masked and pruned models achieved identical MMLU accuracy as the unmodified versions.
This confirms that the selected heads are sufficient to drive retrieval and that the G&A mechanism is not unique to Transformers: despite differences in architecture, SSM-based models converge on the same two-stage strategy.
Nevertheless, the attention patterns differ in texture. As shown in Figure 5, attention maps in Mamba-based models are smoother and less sharply localized. This softness limits their ability to isolate segments cleanly:
  • Gather Heads in SSMs often leak attention outside segment boundaries.
  • Aggregate Heads struggle to sharply focus on the correct summary token.
These imperfections stem from the continuous nature of SSM hidden states, which blur boundary distinctions. We explore these architectural constraints further in Section 5.

### v2

@@ 3 A Case Study: Retrieval in MMLU
The MMLU benchmark (Hendrycks et al., 2021) spans 57 diverse tasks and is often viewed as a test of world knowledge. However, its multiple-choice format also demands retrieval, requiring models to map question-answer context to the correct letter label (A–D), as illustrated below:
@@ D. Access Port
Despite matching Transformers performance on many knowledge benchmarks, SSMs underperform on MMLU unless trained substantially longer (Waleffe et al., 2024). Given their known retrieval limitations (Jelassi et al., 2024), we ask whether this gap stems from a failure of knowledge or retrieval. To investigate, we analyze three representative models: Llama-3.1-8B-Instruct (Transformer), Falcon-Mamba-7B-Instruct (Mamba-1), and Llamba-8B-Untied (Mamba-2). We compare their performance on MMLU versus benchmarks that emphasize factual reasoning with minimal reliance on in-context retrieval. For clarity, we omit instructional and embedding-tying suffixes from model names in our discussion.
Our investigation reveals two key findings:
@@
    Models can perform well on MMLU with limited knowledge, suggesting that the task primarily reflects in-context retrieval for the evaluated models. This adds to concerns raised by prior work (Wang et al., 2024b; Zhao et al., 2024).
@@
These insights, combined with earlier findings about perplexity gaps between attention-based models and SSM variants (Arora et al., 2023), support our main claim: the general performance gap between Transformer- and SSM-based language models can be traced to specific algorithmic capabilities within a small subset of heads.
Our analysis proceeds bottom-up: we begin by contrasting MMLU with these knowledge-focused tasks (Section 3.1), then localize MMLU performance to a single critical layer (Section 3.2). We show coordination between two layers (Section 3.3), narrow the effect to a small set of heads (Section 3.4), and demonstrate that these implement a shared retrieval mechanism across models (Section 3.5).
@@
Throughout the paper, we evaluate performance on two categories of benchmarks: MMLU and what we term Knowledge-Focused Tasks. MMLU’s format distinctively presents multiple-choice questions, requiring models to process lengthy contexts, then identify the correct option by outputting the corresponding letter label (A, B, C, or D). While MMLU is widely used, studies have noted that strong performance may not reflect true understanding or knowledge (Alzahrani et al., 2024; Wang et al., 2024b; Zhao et al., 2024).
In contrast, Knowledge-Focused Tasks include ARC-Challenge and ARC-Easy (Clark et al., 2018), PIQA (Bisk et al., 2019), Winogrande (Sakaguchi et al., 2019), OpenBookQA (Mihaylov et al., 2018), and HellaSwag (Zellers et al., 2019). These benchmarks primarily assess factual knowledge and commonsense reasoning with minimal reliance on recent context retrieval. The core difference lies in how predictions are evaluated: MMLU requires selecting a letter label tied to the correct answer, while the other benchmarks score the likelihood of complete answer texts—shifting the focus from label retrieval to how well the model’s internal knowledge supports each option.
@@ 3.2 A Crucial Layer for MMLU
Motivated by the struggle of SSMs with MMLU (Waleffe et al., 2024) and by the fact that Transformers use specialized components to solve multiple-choice questions (Lieberum et al., 2023), we examine whether solving this task utilizes different components in SSM-based language models as well. Since all language models utilize residual skip connections, eliminating a layer could reveal its specific contribution while keeping the rest of the information flowing.
We iteratively remove layers from the end of the model and evaluate performance on both MMLU and the knowledge-focused tasks. Figure 2 presents performance trends for Llama-3.1-8B as layers are progressively removed, and Figure 8 extends this analysis to Falcon-Mamba-7B and Llamba-8B. For Llama-3.1-8B and Llamba-8B, we ablate both the mixer and its downstream MLP, since the latter likely depends on the former. However, Figure 9 shows that similar trends hold even when only the mixer is removed.
All architectures reveal a striking pattern: while performance on knowledge-focused benchmarks declines steadily as layers are pruned, MMLU performance remains stable—until a single layer is removed, at which point it drops sharply. This suggests that a specific layer encodes the critical skills required for MMLU. High MMLU scores can persist even when most knowledge capacity is lost, highlighting the role of targeted retrieval components over general knowledge.
Figure 3: Performance of Llama-3.1-8B-Minimal with layers removed and restored. We remove one layer at a time, evaluate performance on MMLU and knowledge tasks’ score, and then restore the layer before proceeding to the next. The results show that both L16 and L17 are critical for MMLU performance. Pruning either of these final layers causes a large drop in MMLU accuracy, while scores on knowledge tasks decline only slightly.
@@ 3.3 MMLU Relies on Two Crucial Layers
Following Figure 2, we define the minimal model as the smallest subnetwork that includes all layers up to the one before the MMLU performance drop. We now repeat the previous experiment, this time on the minimal model, with a slight modification. Instead of cumulatively removing one layer at a time from the end of the network, we remove a single layer, evaluate the model performance without it, and then restore the layer before proceeding to the next.
Figure 3 illustrates the results of this experiment in Llama-3.1-8B, and Figure 10 shows the results for Falcon-Mamba-7B and Llamba-8B as well. The findings reveal a surprising insight on the minimal model: not one, but two layers are critical for the performance of MMLU. Unlike earlier layers, the last two layers of the minimal model show no correlation between MMLU and knowledge task performance. Pruning either layer significantly drops MMLU accuracy while minimally affecting knowledge-focused tasks, underscoring their essential role.
@@
Table 1: Evaluating the Role of Critical Heads in Llama-3.1-8B-Minimal’s Performance. This table presents the impact of two critical attention heads (L16H22 and L17H24) on MMLU performance while keeping all preceding layers intact. Four configurations were tested: (1) both heads retained, (2) only the last-layer head retained, (3) only the second-to-last-layer head retained, and (4) neither head retained. Results show that these heads must work together to sustain high MMLU performance ($>$66%), while removing either or both drops accuracy to random guessing (25%).
@@ 3.4 Zooming In: MMLU Relies on Two Crucial Heads
Having established the importance of two specific layers in Section 3.3, we now refine our analysis from the layer level to the head level, investigating how individual attention heads within these layers contribute to performance.
To identify crucial heads, we conduct ablation studies on the minimal model defined in Section 3.3, systematically zeroing out individual head outputs in the two critical layers and measuring the impact on both MMLU and knowledge-focused tasks. This analysis reveals a remarkably small subset of heads whose removal drastically reduces MMLU accuracy while keeping knowledge benchmark performance intact. We designate these as “critical heads” — specifically, the heads whose removal significantly affects MMLU scores.
To further validate the importance of these critical heads, we remove all noncritical heads from the last two layers and evaluate performance across four configurations: (1) no critical heads retained, (2) only critical heads from the last layer retained, (3) only critical heads from the second-to-last layer retained, and (4) critical heads in both layers retained. The results, presented in Table 1 for Llama-3.1-8B and in Table 4 for the other models, demonstrate that MMLU performance depends on just a few mixing heads rather than the model’s broader knowledge representation. Disabling these specific heads causes accuracy to plummet to random guessing levels. Remarkably, in Llama-3.1-8B, a single head is sufficient to achieve over 66% MMLU accuracy, while in Falcon-Mamba-7B, just four Mamba-1 channels reach 52.26% accuracy.
@@ 3.5 Retrieval Emerges as a Key Factor in MMLU
Earlier results showed that two attention heads from separate layers work together to solve MMLU across architectures. We now provide evidence that these heads form part of a broader in-context retrieval mechanism, validated through a synthetic key-value (KV) retrieval task.
In this task, the model is given a dictionary and must recall the value of a queried key, for example,
@@
We tested 55 configurations with growing numbers of key-value pairs, each using 1,000 synthetic samples. For each configuration, we evaluated Llama-3.1-8B in three settings: intact, with L17H24 removed, and with L16H22 removed (the heads identified as critical in Section 3.4). The full evaluation setup and prompt format are discussed in Section A.5.
As shown in Figure 4, removing either L16H22 or L17H24 significantly impairs retrieval, confirming their central role. Notably, the impact of removing L17H24 increases with dictionary size, suggesting its contribution grows with task complexity. These findings extend previous research: the retrieval heads we identified align with and generalize the Content Gatherer head described by Lieberum et al. (2023) and the retrieval mechanisms documented by Wu et al. (2024). Importantly, our work demonstrates that similar retrieval mechanisms also emerge within SSM-based language models. The strong link between retrieval and MMLU performance suggests that the Transformer–SSM gap may stem from differences in a small number of retrieval-critical heads rather than fundamental architectural limitations. This reframes the performance gap as a targeted functional difference—where SSMs may lack specific retrieval capabilities—rather than a broad modeling inadequacy.
@@ 4 Gather-and-Aggregate Heads
We now examine how the critical retrieval heads identified in Section 3 across architectures implement the same Gather-and-Aggregate (G&A) mechanism. The mechanism unfolds in two stages. First, a Gather Head isolates meaningful segments within the input—such as multiple-choice answers—and summarizes each into a representative token. Then, a later Aggregate Head uses these summaries to retrieve the information most relevant to the prompt.
While previous work has described retrieval-oriented heads in Transformers—such as the “Correct Letter” heads identified by Lieberum et al. (2023)—these analyses focused narrowly on multiple-choice tasks. Our findings show that such heads are part of a broader retrieval mechanism that generalizes beyond this format. This also refines the interpretation from Wu et al. (2024), who grouped retrieval heads under a single role; in contrast, we find they specialize in complementary functions within a coordinated process. Finally, we show that this mechanism also emerges in SSM-based models, despite their lack of explicit attention. These models exhibit smoother and less discrete patterns, but still learn to implement a G&A strategy using different architectural tools.
This section unpacks the mechanism in detail. We begin with how it operates in Transformers (Sections 4.1 and 4.2), then present experimental evidence that it also underlies retrieval in SSMs (Section 4.3).
@@ 4.1 Gather Heads in Transformers
A Gather Head compresses semantically related tokens within a segment into a single representative token. In Transformer models, this typically occurs through attention patterns that cause the final token of each segment to “attend to” all prior tokens in the same segment, effectively summarizing them into a compact representation (see Figure 1(a)). For example, consider the multiple-choice format introduced in Section 3:
@@
While effective, this process is not always clean: attention sometimes leaks to unrelated tokens—particularly “attention sinks” Xiao et al. (2024) used to stabilize processing over long contexts.
@@ 4.2 Aggregate Heads in Transformers
Once the Gather Head condenses segments into representative tokens, an Aggregate Head—in a subsequent layer—retrieves and combines these summaries to inform the model’s prediction. This head typically attends from a target token (e.g., the final “Answer:” position) over all gathered summaries, assigning greater weights to the most relevant one (see Figure 1(b)). This yields a sparse linear combination over segment summaries, often resembling an argmax-like operation (Tulchinskii et al., 2024):
@@ 4.3 G&A in Transformers and SSMs: Experimental Evidence
The G&A pattern is well-documented in Transformers—but do SSM models like Mamba implement the same structure? We test this via a two-part experiment that isolates and evaluates specific heads: (1) Pruning: We prune each model to a minimal configuration (as described in Section 3.3), zeroing out all heads in the final two layers except those suspected to implement G&A, as illustrated in Table 4. (2) Masking: At runtime, we constrain the preserved heads’ attention patterns: Gather Heads may attend only within each segment and to its final token, and Aggregate Heads may attend only to those final summary tokens (see Figure 6).
Across all models—Llama-3.1-8B (Transformer), Llamba-8B (Mamba-2), and Falcon-Mamba-7B (Mamba-1)—those pruned and masked to retain only G&A heads achieve the same MMLU accuracy as their unmodified counterparts. This confirms that the selected heads are sufficient to drive retrieval and that the G&A mechanism is not unique to Transformers: despite differences in architecture, SSM-based models converge on the same two-stage strategy.
Nevertheless, the attention patterns differ in texture. As shown in Figure 5, attention maps in Mamba-based models are smoother and less sharply localized. This softness limits their ability to isolate segments cleanly: Gather Heads in SSMs often leak attention outside segment boundaries, and Aggregate Heads struggle to sharply focus on the correct summary token. These imperfections stem from the continuous nature of SSM hidden states, which blur boundary distinctions. We explore these architectural constraints further in Section 5.